OpenStudy (anonymous):

this question is nasty... would prefer ANY help offered (lemme type in comments)

OpenStudy (anonymous):

h(x)=\[1\div5^{2\cos (x-\Pi \div2)+1}\] find 2 functions, f(x) and g(x) such that \[\left( f o g \right)(x)=h(x)\]

terenzreignz (terenzreignz):

haha, I feel like being a smart*ss and letting f(x) = x g(x) = h(x), but I'm guessing that's not what you want? :P

OpenStudy (anonymous):

haha, after a few attempts, that ones better than my answer :P i literally have no idea where to start with this one. (i wrote bacon)

terenzreignz (terenzreignz):

How about f(x) = 1/(5^x) g(x) =2cos(x - *pi* / 2) + 1

terenzreignz (terenzreignz):

there are literally infinitely many possible answers, I'm surprised your instruction wasn't more specific

OpenStudy (anonymous):

i wrote the question down word for word o.O if you want more insight to what its asking: Its basically looking for two functions that when multiplied together give you that ex-wife looking formula up there... f(x)=1 g(x)=h(x) problem math?

terenzreignz (terenzreignz):

Please keep in mind that this strictly isn't 'multiplication' of functions, but 'composition' also, if f(x) = 1 and g(x) = h(x), then f o g (x) = f(g(x)) = 1 lol, I think you meant f(x) = x